Circular No. 6941

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE NOVA IN OPHIUCHUS S. Nakano, Sugano, Japan, reports the discovery of a possible nova

(mag 9.5) by Kesao Takamizawa, Saku-machi, Nagano-ken, which was found on two T-Max 400 films (each exposure ∼ 4.5 min) taken on June 15.561 UT with a 0.10-m f/4.0 lens. Takamizawa reports the position of the possible nova as α = 17^{h}32^{m}00^{s}.0, δ = −19^{o}13′56′′ (equinox 2000.0). No star is present at this location on 48 previous patrol films taken by Takamizawa between 1994 Feb. 17 and 1998 May 19. R. and Y. Kushida report the following position end figures for the variable star from a poor CCD image obtained on June 16.531, when they estimated V = 10.0: 31^{m}59^{s}.82, 57′′.0 (uncertainty stated as 1′′ or more; average of three frames). There is a red USNO A1.0 star (B = 18.0, R = 16.8) at position end figures 59^{s}.79, 55′′.0. T. Kato, Kyoto University, notes that the possible nova lies within the 16′′ error circle of the ROSAT source 1RXS 173200.0–191349.

1998 KY_{26}

P. Pravec and L. Šarounová, Ondřejov Observatory, report: “Our photometric observations of this Apollo-type object from Ondřejov during June 1.977–2.040 and 2.875–3.037 UT revealed a synodic period of 10.702 ± 0.003 min; periods other than half-integer multiples of this period are apparently ruled out. The lightcurve has most of its signal in the second harmonic (i.e., it shows two maxima/minima pairs per period), and the peak-to-peak amplitude is 0.3 mag. We interpret the period as being the rotation period of the asteroid. This result is in agreement with the radar observations by Ostro et al. (IAUC 6935). This rotation period indicates that the object is a monolithic body. A body composed of pieces held together by self-gravitation only (the so-called ‘rubble pile’ structure; cf. Harris 1996, Lunar Planet. Sci. 27, 493) would require an unrealistic bulk density of more than 340 g/cm^{3}. Our composite lightcurve is presented at

http://sunkl.asu.cas.cz/~ppravec/98ky26.htm.”

COMET C/1997 J2 (MEUNIER-DUPOUY) Visual m_{1} estimates: May 1.04 UT, 11.0 (M. Meyer, Frauenstein, Ger-

many, 0.25-m reflector); 18.00, 11.4 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); June 3.98, 10.7 (M. Reszelski, Szamotuly, Poland, 0.25-m refl.); 13.89, 11.2 (A. Baransky, Kiev, Ukraine, 0.20-m reflector).

1998 June 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green